OpenStudy (anonymous):

write each equation in the system of equations in slope intercept form. use the slope inter. form to determine the number of solutions the system had. {6x-7y+5....... -12x+14y+-20... A. infinite number of solutions. b. one solutions. c. no solution. c. cannot be determined....

OpenStudy (anonymous):

the = signs are missing..

OpenStudy (anonymous):

yup proper sign

OpenStudy (anonymous):

omg.. im sorry.. its 6x-7y=5 and -12x+14=-20

OpenStudy (anonymous):

still i suppose ans is no solution

OpenStudy (anonymous):

they represent two distinct parallel lines hence no solution

OpenStudy (anonymous):

thx... but how did you figure? because i have a few more questions like that, and i want to be able to answer properly. :)

OpenStudy (anonymous):

x-3y+5... -5x+15=-25 is another one

OpenStudy (anonymous):

again improperly stated

OpenStudy (anonymous):

for eq ax+by=c and px+qy=r if a/p <>b/q then lines are not parallel and unique solution ifa/p=b/q =c/r then same line and infinite number of solutions if a/p=b/q but not equal toc/r then distinct parallel lines and hence there is no solution

OpenStudy (anonymous):

i hope it will help u to do it on ur own

OpenStudy (anonymous):

you're awesome.. thanks! :)

OpenStudy (anonymous):

welcome dear